OpenStudy (anonymous):

Write (-2,-9), (-1,7) in slope-intercept form, standard form, and point slope form.

OpenStudy (anonymous):

y-7 = 16(x+1)

OpenStudy (anonymous):

y=16x + 23

OpenStudy (anonymous):

which is in d clope intercept form

OpenStudy (anonymous):

slope*

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

first calculate the slope of the line by using the two points given(m). Now use this m. First for slope point form. we know that (y-y1)=m(x-x1) Now put the value of the first point that you have selected and solve it. 2. Once you fine the eqn in terms of x and y, take y to one side of equality and all the other term on the other side. divide the eqn by the coefficient of y. you get the slope intercept form. the standard for is nothing but the general case of point slope form. Instead of calculating the slope of the question now replace the entire expression of slope calculation in place of m and the rest of the equation remains the same...